OpenStudy (anonymous):

Using Power reduction to solve cos^4t dt?

OpenStudy (isaiah.feynman):

Invalid question.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

invalid?

OpenStudy (anonymous):

@Isaiah.Feynman its cos^4 (t)

OpenStudy (isaiah.feynman):

So what do you want to solve there?

OpenStudy (anonymous):

well we have to use power reduction. wouldnt you reduce it to cos^2(t) cos^2(t)?

OpenStudy (isaiah.feynman):

Yeah.

OpenStudy (anonymous):

then(1+cos2x/2)(1+cos2x/2) ?

OpenStudy (isaiah.feynman):

Where did you get that from?

OpenStudy (anonymous):

cos^2x dx= integral of 1+cos2x/2

OpenStudy (isaiah.feynman):

No. What you mean is the trig identity \[\cos^{2}x = 1-\sin^{2}x\]

OpenStudy (anonymous):

its a power reduction formula?

OpenStudy (isaiah.feynman):

No.

zepdrix (zepdrix):

You could use the ummmm Half-Angle Formula's a bunch of times. Or you could use the Cosine Redux Formula:\[\Large I_n\quad=\quad \frac{1}{n}\cos^{n-1}x \sin x +\frac{n-1}{n}I_{n-2}\]

zepdrix (zepdrix):

Where I_n is,\[\Large I_n\quad=\quad \int\limits \cos^n x\;dx\]

zepdrix (zepdrix):

Although, if you haven't learned about that yet, then maybe half-angles are easier :)

zepdrix (zepdrix):

\[\Large \cos^4x \quad=\quad [\cos^2x]^2 \quad=\quad \left[\frac{1}{2}(1+\cos2x)\right]^2\]

zepdrix (zepdrix):

Gotta expand that out and do some more half angle after that ^^

OpenStudy (anonymous):

1/4(x+1/2sin2x)?

zepdrix (zepdrix):

Hmm I'm confused.. did you ignore the square on the outside? D:

OpenStudy (anonymous):

wouldnt it just be another 1/2?

zepdrix (zepdrix):

No :(

OpenStudy (anonymous):

hahaha. sorry. I dont really understand half angles

zepdrix (zepdrix):

This is what expanding the brackets gives us,\[\Large \frac{1}{4}(1+2\cos2x+\cos^22x)\]That part make sense? :o

zepdrix (zepdrix):

I multiplied out the outer square

OpenStudy (anonymous):

ohhh i know this. duh

zepdrix (zepdrix):

What is that reindeer doing on that lady's face? +_+ That's kinda rude.

zepdrix (zepdrix):

You'll want to apply the Half-Angle Formula again to this orange term before you can integrate. \[\Large \frac{1}{4}(1+2\cos2x+\color{orangered}{\cos^22x})\]

OpenStudy (anonymous):

Doing reindeer things :D

zepdrix (zepdrix):

pshhh

OpenStudy (anonymous):

hmm so youd plug in the 1+cos2x/2 into it?

zepdrix (zepdrix):

No the angle `doubles` each time we apply the rule. So we multiply another 2 inside of the cosine.

zepdrix (zepdrix):

\[\Large \cos^22x\quad=\quad \frac{1}{2}(1+\cos(2\cdot2x)) \quad=\quad\frac{1}{2}(1+\cos4x)\] Understand how that works? :) Dubs. The. Angle.

OpenStudy (anonymous):

yup! thanks<3

OpenStudy (isaiah.feynman):

@skay your smartscore graph is a straight line. Can you get it's equation? :P